# Hilbert-Kunz multiplicity of fiber products and idealization rings

### v3

@@ Abstract.
The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings. We calculate the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, with $R$, $S$ and $T$ being Noetherian local rings with common characteristic and residue field. Inspired by a work of Strazzanti and Armengou we calculate the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. We explore certain examples and structural consequences and obtain some bounds for the Hilbert-Kunz multiplicity.
@@
Hilbert-Kunz, multiplicity, fiber product, idealization
@@
13D40, 13H15, 13H05
Corresponding author: P. D. C. Martins
The first author was supported by grant 2019/21181-0, São Paulo Research Foundation (FAPESP). The second author was supported by grant 2022/12114-0, São Paulo Research Foundation (FAPESP)
@@ 1. Introduction
In 1983, Monsky [18] introduced a refined version of multiplicity specific to Noetherian local rings of prime characteristic, termed the Hilbert-Kunz multiplicity. It’s worth noting that this notion was also introduced by Kunz [16] in 1969, but Monsky provided a more explicit definition. While this invariant remains largely enigmatic, its study has revealed intriguing connections with the classical multiplicity, albeit often exhibiting more intricate behavior. Unlike the Hilbert-Samuel multiplicity, which is always an integer, the Hilbert-Kunz multiplicity may not be. However, this characteristic enables us to regard it as a more nuanced invariant. Similar to the classical multiplicity, it is capable of discerning regular rings. The study of Hilbert-Kunz multiplicity has engaged numerous researchers, as evidenced by works such as [18], [24], [8], and others.
@@
Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finite $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
@@
where $q=p^{e}$ and $I^{[q]}$ is the ideal generated by $x^{q}$, $x\in I$. If $M$ is a finite $R$-module $M/I^{[q]}M$ has finite lenght. For simplicity, we can put $e_{HK}(I):=e_{HK}(I,R)$ and $e_{HK}(R):=e_{HK}(\mathfrak{m})$, if it doesn’t cause confusion. Notice that the limit of the right-hand side of (1.1) always exists and is a real number; see, e.g., Monsky [18].
In [17], Kunz observed that the number $\ell_{e}(R):=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$. In some sense, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$.
If $R$ is regular, then $e_{HK}(I,R)=\ell_{R}(R/I)$ for every $\mathfrak{m}$-primary ideal $I$ of $R$ and $e_{HK}(R)=1$. Thus, it is natural to ask: ”If $e_{HK}(R)=1$, is $R$ then regular?”. There is a positive answer to the above question, given by Watanabe and Yoshida [23], when $R$ is unmixed, i. e., given an unmixed local ring $(R,\mathfrak{m})$ of characteristic $p>0$, if $e_{HK}(R)=1$, then $R$ is regular. A somewhat simpler proof can be found in [15]. Watanabe and Yoshida [24] also proved that it holds when $R$ is a two-dimensional Cohen-Macaulay local ring.
The main purpose of this paper is to provide new formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and consequently produce criteria regarding the regularity of such rings. The study of fiber products and idealization rings has recently received special attention (see for instance [2, 7, 12, 13, 19, 20]). One reason to obtain formulas for the Hilbert-Kunz multiplicity of fiber product rings comes from the fact that the structure of the fiber product ring changes severely, depending on the rings considered (see Corollary 3.6 and Proposition 3.7).
Let us now briefly describe the contents of the paper. Section 2 establishes the main definitions and facts for the rest of the paper. Section 3 is devoted to study the Hilbert-Kunz multiplicity of fiber product rings. Some structural results and examples are also given. Motivated by a result of [21], whose obtain as a consequence a formula for the Hilbert-Kunz multiplicity for idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$, in the Section 4 we generalize and study the Hilbert-Kunz multiplicity of the idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. Some structural results are provided. In the last section, we discuss two conjectures concerning lower bounds for the Hilbert-Kunz multiplicity. In particular, we provide new classes of $d$-dimensional Noetherian local rings of characteristic $p$ whose satisfy a conjecture posed by Watanabe and Yoshida [25], that asks if $e_{HK}(R)$ is bounded by $e_{HK}(R_{p,d})$, where $(R,\mathfrak{m},k)$ is a $d$-dimensional unmixed local ring of characteristic $p>2$ and dimension $d\geq 2$ and $R_{p,d}$ is the hypersurface $k[[X_{0},X_{1},\ldots,X_{d}]]/(X_{0}^{2}+\cdots+X_{d}^{2})$, with $k=\overline{\mathbb{F}_{p}}$.
@@ 2. Setup and Basic facts
We will start by recalling some fundamental properties of the Hilbert-Kunz multiplicity. After that, we will introduce the notion of fiber products and idealization rings to study their Hilbert-Kunz multiplicities. We observe that these two constructions work in general, when the ring characteristic is not necessarily positive.
Recall that if $R$ is a Noetherian local ring and $M$ is a finite $R$-module, then $\dim M$ is defined as $\dim M:=\dim(R/\operatorname{Ann}_{R}(M))$. Also, if $R$ is a local domain, then the rank of $M$ is $\operatorname{rk}_{R}M:=\dim_{Q}(Q\otimes_{R}M)$, where $Q$ is the total quotient ring of $R$.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$. Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finite $R$-module. Then the following statements hold:
@@
    (Additivity in short exact sequences) If $0\longrightarrow L\longrightarrow M\longrightarrow N\longrightarrow 0$ is a short exact sequence of finite $R$-modules, then
@@
    where $\operatorname{Assh}(R)$ denotes the set of prime ideals $\mathfrak{p}$ of $R$ with $\dim R/\mathfrak{p}=\dim R$. In particular, if $R$ is a local domain and $M$ is a torsion-free $R$module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(R)$
@@
    If $x\in I$ is $R$-regular, then $e_{HK}(I)\leq e_{HK}(I/xR)$.
  • (9)
    (Localization) $e_{HK}(R_{\mathfrak{p}})\leq e_{HK}(R)$ for any prime ideal $\mathfrak{p}$ such that $\dim R/\mathfrak{p}+\operatorname{height}\mathfrak{p}=\dim R$
  • (10)
    If $(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
Fiber product rings
@@
is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [2, Lemma 1.2]). Let $\eta_{R}:R\times_{T}S\twoheadrightarrow R$ and $\eta_{S}:R\times_{T}S\twoheadrightarrow S$ be the natural projections $(r,s)\mapsto r$ and $(r,s)\mapsto s$, respectively. Then $R\times_{T}S$ is represented as a pullback diagram:
@@
In general, when we use the symbols $R$, $S$ or $T$, we assume that $R$, $S$ and $T$ are non-zero Noetherian local rings with the maximal ideal $\mathfrak{m}_{R}$, $\mathfrak{m}_{S}$ and $\mathfrak{m}_{T}$, respectively, and with $k$ being the common residue field. Remark that each (finite) module over $R$, $S$ or $T$ acquires a canonical stucture of (finite) $R\times_{T}S$-module via Diagram (2.5).
We will state three already-known facts about fiber product rings to be used in Section 3.
@@
There is the exact short sequence of $R\times_{T}S$-modules
@@
Idealization of a module
Next construction was introduced by Nagata and its nice properties were studied by Anderson and Winders [3].
@@
Taking a finite $R$-module $M$ over a $d$-dimensional Noetherian local ring of prime chatecteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$, it is well known that the idealization ring $R\ltimes M$ is a $d$-dimensional Noetherian local ring of positive chatecteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field k. Also, we can consider $R\hookrightarrow R\ltimes M$ by sending $r\mapsto(r,0)$. Then $R\ltimes M$ is a local $R$-algebra with the same residue field of R. It is discussed in [3].
@@ 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2. Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively. In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$. During this section, fiber product rings $R\times_{T}S$ will be considered under these conditions. To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
@@
Let $R\times_{T}S$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S>\dim T.\end{array}\right.$
@@
Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.4 and the additivity of Hilbert-Kunz multiplicity in exact short sequences, we get
@@
Applying the same argument to the exact sequence
we obtain $e(I,R\oplus S)=e(I,R)+e(I,S)$ and then
$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T)$
Note that the dimensions of $R$, $S$ and $T$ as $R\times_{T}S$-modules coincide with their respective Krull dimensions. Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.5 and item (3) of Proposition 2.1 we get the desired formulas. ∎
Next proposition provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
3.2 Proposition.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). Then the following equality holds
$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S>\dim T.\end{array}\right.$
@@
As an immediate consequence of Lemma 3.1, we obtain the formulas
(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S>\dim T.\end{array}\right.$
Now, we will discuss why we can replace $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$ and $e_{HK}(\mathfrak{m},T)$ in (3.4) by $e_{HK}(R)$, $e_{HK}(S)$ and $e_{HK}(T)$, respectively. Note that $R\times_{T}S$ also has characteristic $p$. In general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomophism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so it is easy to see that $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$. Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module $N$, we obtain:
@@
Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas, by observing that we need only to consider the components of the fiber product $R\times_{T}S$ whose have dimension equal to $\dim R\times_{T}S$. ∎
3.1 Importance of the formula.
Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular. Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [25, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings. However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed. In this sense, by using existing results, Proposition 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring. Below, we provide one example among many to illustrate this fact:

### v4

# Hilbert-Kunz Multiplicity of Fiber Product Rings and Nagata Idealizations

@@ Abstract.
The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations. We give explicit formulas for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, where $R$, $S$, and $T$ are Noetherian local rings sharing the same characteristic and residue field. We compute the Hilbert-Kunz multiplicity of a general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module. Additionally, we provide examples, structural results and establish new bounds for the Hilbert-Kunz multiplicity.
@@
Fiber product ring, Nagata idealization, Hilbert-Kunz multiplicity, Amalgamated duplication
@@
13H15, 13H10
Corresponding author: Paulo Martins
@@ 1. Introduction
In 1983, Monsky [23] introduced a refined version of multiplicity specific to Noetherian local rings of prime characteristic, termed the Hilbert-Kunz multiplicity. It’s worth noting that this notion was also introduced by Kunz [21] in 1969, but Monsky provided a more explicit definition. While this invariant remains largely enigmatic, its study has revealed intriguing connections with the classical multiplicity, albeit often exhibiting more intricate behavior. Unlike the Hilbert-Samuel multiplicity, which is always an integer, the Hilbert-Kunz multiplicity may not be. However, this characteristic enables us to regard it as a more nuanced invariant. Similar to the classical multiplicity, it is capable of discerning regular rings.
@@
Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finitely generated $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
@@
where $q=p^{e}$ and $I^{[q]}$ is the ideal generated by $x^{q}$, $x\in I$. If $M$ is a finitely generated $R$-module $M/I^{[q]}M$ has finite lenght. For simplicity, we can put $e_{HK}(I):=e_{HK}(I,R)$ and $e_{HK}(R):=e_{HK}(\mathfrak{m})$, if it doesn’t cause confusion. Notice that the limit of the right-hand side of (1.1) always exists and is a real number; see, e.g., Monsky [23].
In [22], Kunz observed that the number $\ell_{e}(R)=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$. If $R$ is a regular local ring, then $e_{HK}(R)=1$. Otherwise, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$. The study of fiber products and idealization rings has recently garnered significant attention (see, for instance, [4, 3, 15, 16, 17, 18, 24, 26]). Freitas and Perez [18] obtained explicit formulas for the Hilbert multiplicity of graded fiber product rings and Strazzanti and Armengou [27] obtained, as a consequence of their results for some quadratic quotients of the Rees algebra, a formula for the Hilbert-Kunz multiplicity of Nagata idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$. These results motivated us to obtain explicit formulas for the Hilbert-Kunz multiplicity of a fiber product ring $R\times_{T}S$ (Theorem 3.2) and for a general Nagata idealization $R\ltimes M$ (Corollary 4.3). These formulas have yielded new results on the structure of fiber product rings, as well as sharper bounds for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations.
Let us briefly describe the contents of the paper. Section 2 establishes our terminology and some preparatory facts that will be needed throughout the paper. Motivated by the formulas of Freitas and Perez [18] for the Hilbert multiplicity of graded fiber product rings, in Section 3, we obtain explicit formulas for the Hilbert-Kunz multiplicity of a fiber product ring. We provide a simpler proof of a result established by Freitas and Lima [15] that a non-trivial general fiber product ring $R\times_{T}S$ is not regular. Additionally, we present a formula for the Hilbert-Kunz multiplicity of the amalgamated duplication $R\bowtie I$ and examples. In Section 4, the main focus is to obtain an explicit formula for the Hilbert-Kunz multiplicity of the general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module. Section 5 is devoted to studying the behavior of singularities of fiber product rings using a Hilbert-Kunz multiplicity approach. In the final section, we establish bounds for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations. Furthermore, we extend the formula for the Hilbert-Kunz multiplicity of fiber product rings to the more general case of multi-factor fiber products. As an application of our results, we identify new classes of singular rings that satisfy the celebrated Watanabe-Yoshida conjecture.
@@ 2. Setup and Basic Facts
We begin by recording some fundamental properties of Hilbert-Kunz multiplicity. We then introduce the definitions of fiber products and idealization rings, along with relevant facts that will be used to explore their Hilbert-Kunz multiplicities in the next sections.
@@
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$. Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finitely generated $R$-module. Then the following statements hold:
@@
    (Additivity in short exact sequences) If $0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0$ is a short exact sequence of finitely generated $R$-modules, then
@@
    where $\operatorname{Assh}(R)$ denotes the set of prime ideals $\mathfrak{p}$ of $R$ with $\dim R/\mathfrak{p}=\dim R$. In particular, if $R$ is a local domain and $M$ is a torsion-free $R$-module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(I,R)$ .
@@
    If $f:(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
The study of lower bounds for the Hilbert-Kunz multiplicity has been explored by several authors (see e.g., [29, 30, 31, 8, 1]). Using an elegant argument, Aberbach and Enescu [1] resolved the fundamental problem of obtaining bounds that are independent of the characteristic in all dimensions.
We say that a local ring $(A,\mathfrak{m}$) is unmixed (resp. quasi-unmixed) if $\dim\widehat{A}/\mathfrak{p}=\dim\widehat{A}$ for every associated (resp. minimal) prime ideal of $\widehat{A}$. Here, $\widehat{A}$ denotes the $\mathfrak{m}$-adic completion of $A$.
2.2 Theorem (Aberbach-Enescu).
Let $(R,\mathfrak{m},k)$ be an unmixed ring of characteristic $p$ and dimension $d\geq 2$. If $R$ is not regular, then
$e_{HK}(R)\geq 1+\frac{1}{d(d!(d-1)+1)^{d}}.$
2.1. Fiber product rings
@@
is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [4, Lemma 1.2]). Let $\eta_{R}:R\times_{T}S\twoheadrightarrow R$ and $\eta_{S}:R\times_{T}S\twoheadrightarrow S$ be the natural projections $(r,s)\mapsto r$ and $(r,s)\mapsto s$, respectively. Then $R\times_{T}S$ is represented as a pullback diagram:
@@
In general, when we use the symbols $R$, $S$ or $T$, we assume that $R$, $S$ and $T$ are non-zero Noetherian local rings with the maximal ideal $\mathfrak{m}_{R}$, $\mathfrak{m}_{S}$ and $\mathfrak{m}_{T}$, respectively, and with $k$ being the common residue field. The fiber product ring $R\times_{T}S$ is non-trivial, i.e., $R\neq T\neq S$, and their maximal ideal is denoted by $\mathfrak{m}:=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}\mathfrak{m}_{S}$. Note that each (finitely generated) module over $R$, $S$ or $T$ acquires a canonical structure of (finitely generated) $R\times_{T}S$-module via Diagram (2.5).
We will state known facts about fiber product rings to be used in the next sections.
@@
There is the short exact sequence of $R\times_{T}S$-modules
$0\rightarrow R\times_{T}S\rightarrow R\oplus S\rightarrow T\rightarrow 0$
@@
2.8 Fact.
([26, Fact 2.9]) Let $P=R\times_{k}S$ be a non-trivial fiber product. Then
$\operatorname{Spec}P=\{\mathfrak{m}_{R}\oplus\mathfrak{q}\mid\mathfrak{q}\in\operatorname{Spec}S\}\cup\{\mathfrak{p}\oplus\mathfrak{m}_{S}\mid\mathfrak{p}\in\operatorname{Spec}R\}.$
Following this notation, we have $\operatorname{ht}(\mathfrak{m}_{R}\oplus\mathfrak{q})=\operatorname{ht}(\mathfrak{q})$ and $P/(\mathfrak{m}_{R}\oplus\mathfrak{q})\cong S/\mathfrak{q}$, so we must then have $\dim(P/(\mathfrak{m}_{R}\oplus\mathfrak{q}))=\dim(S/\mathfrak{q})$ and similarly for the prime $\mathfrak{p}\oplus\mathfrak{m}_{S}$. Futhermore, if $\mathfrak{q}\neq\mathfrak{m}_{S}$, then $P_{\mathfrak{m}_{R}\oplus\mathfrak{q}}\cong S_{\mathfrak{q}}$, and similarly for $\mathfrak{p}\oplus\mathfrak{m}_{S}$.
2.2. Idealization of a module
Next construction was introduced by Nagata, and its nice properties were studied by Anderson and Winders [6].
@@ 2.2. Idealization of a module
Let $M$ be a finitely generated $R$-module over a $d$-dimensional Noetherian local ring $R$ of prime characteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$. It is well known that the idealization ring $R\ltimes M$ is also a $d$-dimensional Noetherian local ring of characteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field $k$. Furthermore, we can embed $R$ into $R\ltimes M$ via $r\mapsto(r,0)$, making $R\ltimes M$ a local $R$-algebra with the same residue field as $R$. This is discussed in detail in [6].
@@ 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the main focus is to present a formula for the Hilbert-Kunz multiplicity of the fiber product ring $R\times_{T}S$, where $R$, $S$, and $T$ are Noetherian local rings with the same characteristic $p>0$ and residue field. In this section, we consider rings $R$, $S$ and $T$ as in Definition 2.3 and adopt the Convention 2.4. To avoid confusion, we note that in this section, $R$, $S$, and $T$ will also be treated as $R\times_{T}S$-modules. Without loss of generality (see Fact 2.6), we may assume that $\dim R\times_{T}S=\dim R\geq\dim S\geq\dim T$.
@@
Let $R\times_{T}S$ be a fiber product ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
@@
Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.5 and the additivity of Hilbert-Kunz multiplicity in short exact sequences, we get
@@
Applying the same argument to the exact sequence $0\rightarrow R\rightarrow R\oplus S\rightarrow S\rightarrow 0$ we obtain $e_{HK}(I,R\oplus S)=e_{HK}(I,R)+e_{HK}(I,S)$ and then
$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T).$
Note that the dimensions of $R$, $S$ and $T$ as $R\times_{T}S$-modules coincide with their respective Krull dimensions. Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.6 and item (3) of Proposition 2.1 we get the desired formulas, nothing that the components of the fiber product that have dimension less than $\dim R\times_{T}S$ have Hilbert-Kunz multiplicity with respect to $I$ equal to zero. ∎
3.2 Theorem.
Let $R$, $S$ e $T$ be Noetherian local rings of characteristic $p>0$. Then the following equality holds
$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
@@
As an immediate consequence of Lemma 3.1, taking $I=\mathfrak{m}$, we obtain the following formulas
(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
Now, we will discuss why we can replace $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$ and $e_{HK}(\mathfrak{m},T)$ in (3.4) by $e_{HK}(R)$, $e_{HK}(S)$ and $e_{HK}(T)$, respectively. It is easy to see that $R\times_{T}S$ also has characteristic $p$. Note that, in general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomorphism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$. Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module NN of finite lenght, we obtain:
@@
Therefore, using the formula for $\dim R\times_{T}S$ from Fact 2.6 together with the surjective ring homomorphisms in diagram (2.5), we may substitute $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$, and $e_{HK}(\mathfrak{m},T)$ in (3.4) with $e_{HK}(R)$, $e_{HK}(S)$, and $e_{HK}(T)$, respectively. Here we observe that it suffices to consider only those components of the fiber product $R\times_{T}S$ whose dimensions are equal to $\dim R\times_{T}S$. ∎
The study of structural properties of fiber product rings has been addressed by several authors, as [17] and [4]. In [25, Proposition 2.6], they proved that a non-trivial fiber product of the form $R\times_{k}S$ is not regular. Freitas and Lima [15, Theorem 4.7(1)] extended this result to any non-trivial fiber product ring using a Betti numbers approach. We give a simpler proof that any non-trivial fiber product $R\times_{T}S$ is not regular. We will make use of the singularness of fiber products in the next sections.
3.3 Theorem.
A non-trivial fiber product $R\times_{T}S$ is not a domain.
Proof.
Since $\pi_{R}$ and $\pi_{S}$ are surjective and not isomorphisms, we have $\ker(\pi_{R})\neq 0$ and $\ker(\pi_{S})\neq 0$. Let $r_{0}\in\ker(\pi_{R})\setminus\{0\}$ and $s_{0}\in\ker(\pi_{S})\setminus\{0\}$. Then,
$\pi_{R}(r_{0})=0=\pi_{S}(0)\Rightarrow(r_{0},0)\in R\times_{T}S,$
$\pi_{R}(0)=0=\pi_{S}(s_{0})\Rightarrow(0,s_{0})\in R\times_{T}S.$
However, $(r_{0},0)\cdot(0,s_{0})=(r_{0}\cdot 0,0\cdot s_{0})=(0,0),$ and both $(r_{0},0)\neq 0$ and $(0,s_{0})\neq 0$. Hence, $R\times_{T}S$ is not a domain. ∎
3.4 Corollary.
Let $R$, $S$, and $T$ be Noetherian local rings. Then the fiber product $R\times_{T}S$ is not regular.
3.1. Amalgamated duplication
Let $R$ be a Noetherian local ring. If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, then $R\times_{R/I}R=R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [11]. Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [4, Theorem 1.8 and 1.9 - 1.9.4].
We can use Theorem 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
3.5 Corollary.
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be a proper ideal of $R$. Then
$e_{HK}(R\bowtie I)=\left\{\begin{array}[]{llll}2e_{HK}(R)-e_{HK}(R/I);&\dim R=\dim R/I\\ 2e_{HK}(R);&\dim R>\dim R/I.\end{array}\right$.
In particular, if $I$ contains a non-zero divisor, then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
Proof.
The proof follows directly using the equality $R\times_{R/I}R=R\bowtie I$ and the formulas given in Theorem 3.2. ∎
3.6 Remark.
The formula given in Corollary 3.5 generalizes a formula obtained in [27, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgamated duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$. In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
Since $R\bowtie I=R\times_{R/I}R$, the next result is a consequence of Theorem 3.3.
3.7 Corollary.
Let $R$ be a Noetherian local ring and let $I$ be a proper ideal of $R$. Then $R\bowtie I$ is not regular.
To end this section, we will use the formulas given in Theorem 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.